# A Physics-Inspired Optimizer: Velocity Regularized Adam

### v1

@@ 5 Benchmarks
We benchmark VRAdam against various optimizers on several datasets for various tasks with diverse architectures. These benchmarks include image classification using a CNN on the CIFAR-10 dataset Krizhevsky and Hinton (2009), language modelling with a Transformer architecture Vaswani et al. (2023) on the WikiText2 dataset Merity et al. (2016), and Generative Flow Networks (GFLowNets) Avidan et al. (2025), on a grid world for sequence generation tasks Chevalier-Boisvert et al. (2023). We also report hyperparameter sensitivity on a diffusion model trained through an UNet architecture Petit et al. (2021) with self attention modules placed in the bottle neck and upper layers for the handwritten digit dataset MNIST LeCun et al. (1998). These benchmarks represent a broad variety of deep learning architectures and tasks, ranging from older techniques for image classification to much newer and actively emerging models and training techniques such as for GFLowNets. We focus on the comparison with AdamW since its exceeding popularity and consistent performance but also report the performance of stochastic gradient descent (SGD) with momentum (Qian, 1999) and root mean square propagation (RMSProp) Ruder (2017). In particular, we demonstrate the excellent convergence properties as well performance improvements in most benchmarks for VRAdam. Note that we run Bayesian optimization for hyperparameter sweeps for the image classification, language modeling and diffusion modeling tasks, with several evaluations in close proximity to the found minimum, and the results are statistically significant. We report the optimal value from this optimization procedure, however, we include shaded error bars in the appendix. The Bayesian optimization for the hyperparameters was conducted with the objective to minimize validation loss, however, we also report test loss. Hyperparameters for the task involving GFlowNets were picked at random due compute constraints. All details regarding model training, setup and datasets along with error envelopes using multiple runs are provided in Appendices B and D.
5.1 Image classification
Image classification serves as a foundational benchmark for deep learning optimizers. This task evaluates VRAdam’s ability to navigate complex, non-convex loss landscapes typical of convolutional neural networks (CNNs) and to achieve good generalization on unseen data. The CIFAR-10 dataset comprises 60,000 32x32 color images categorized into 10 distinct classes, with 6,000 images per class. The dataset is conventionally partitioned into 50,000 images for training and 10,000 for testing. Input images are preprocessed with a normalization scheme based on the per-channel mean and standard deviation of the CIFAR-10 dataset. During training, data augmentation techniques are applied to enhance model robustness and reduce overfitting; these include random horizontal flips and random crops.
The CNN consists of a feature extractor that utilizes three sequential convolutional blocks, where each block contains two 3x3 convolutional layers, each followed by batch normalization and ReLU activation, and concludes with a 2x2 max-pooling operation. These blocks progressively increase the channel depth while reducing the spatial dimensions of the feature maps. The resulting feature maps are then flattened and passed to a classifier module. This classifier comprises a fully-connected layer with a ReLU activation, followed by a final fully-connected layer that outputs logits for the 10 CIFAR-10 classes. Optimizers are tasked with minimizing the Cross-Entropy Loss, a standard loss function for multi-class classification Mao et al. (2023).
Method Validation Loss Test Loss
VRAdam 0.476 0.469
AdamW 0.522 0.565
SGD+Nesterov 0.625 0.620
RMSProp 0.801 0.813
Table 1: Image classification results of optimizers on CIFAR-10 reported through model loss.
5.2 Language modeling
Language modeling is a core task in natural language processing, requiring models to capture complex dependencies and contextual information within sequential data. Here, we assess how well VRAdam compares to other optimizers in training large Transformer models for language modeling. The WikiText-2 dataset, a corpus of over 100 million tokens from Wikipedia articles, is used. The preprocessing step involves data tokenization, and sequence construction.
The Transformer used is a standard encoder-decoder architecture with embedding layers, positional encoding, multi-head attention, and feed-forward networks. The training objective is next-token prediction, where optimizers work to minimize the Cross-Entropy Loss between the model’s predictions and the actual subsequent tokens, with padding tokens ignored during loss calculation. We report NaNs for SGD with momentum and Nesterov, as well as RMSProp due to well-known numerical instability across multiple hyperparameter sweeps Liu et al. (2023b); Vashisht and Ramaswamy (2023); Popel and Bojar (2018).
Method Validation Loss Test Loss
VRAdam 5.99 6.00
AdamW 6.47 6.50
SGD+Nesterov NaN NaN
RMSProp NaN NaN
Table 2: Language modeling results of optimizers on WikiText-2 reported through model loss.
5.3 GFlowNets on GridWorld
This benchmark evaluates the performance of VRAdam on a distinct generative task using GFlowNets. GFlowNets are a class of probabilistic models adept at learning to sample complex, discrete objects $x$ (such as sequences of actions leading to grid patterns) from a target distribution $P^{*}(x)$ that is proportional to a given non-negative reward function $R(x)$, i.e., $P^{*}(x)\propto R(x)$. Unlike traditional reinforcement learning (RL) approaches that often converge to a policy maximizing expected reward (potentially collapsing to a single mode), GFlowNets are designed to learn to sample a diversity of high-reward candidates, making them particularly well-suited for tasks where exploring multiple solutions is crucial. This benchmark tests an ability of an optimizer to handle the unique training objectives of GFlowNets, which involve learning a stochastic forward policy $P_{F}(a_{t}|s_{t};\theta)$ to construct objects step-by-step, and often an auxiliary backward policy $P_{B}(s_{t-1}|s_{t};\phi)$ and a flow function $F(s;\psi)$ over a structured state space $\mathcal{S}$.
[…]
  1. 1.
    A Flow Matching or Trajectory Balance Loss ($\mathcal{L}_{\text{FM}}$): This core component ensures that the learned policy and flow function are internally consistent. This loss component aims to satisfy a detailed balance-like condition, minimizing the squared difference between forward flows $F(s_{t};\psi)P_{F}(a_{t}|s_{t};\theta)$ and backward flows $F(s_{t+1};\psi)P_{B}(s_{t}|s_{t+1};\phi)$ for transitions $(s_{t},a_{t},s_{t+1})$ in sampled trajectories. It also includes a term to match the flow of terminal states $F(s_{T};\psi)$ to their rewards $R(s_{T})$.
  2. 2.
    A Reward Loss Component ($\mathcal{L}_{R}$) and a Entropy Regularization Term ($\mathcal{L}_{H}$) for the policy: These term directly encourage the generation of high-reward states and exploration respectively.
The total loss is a weighted sum: $\mathcal{L}_{\text{GFN}}=w_{\text{FM}}\mathcal{L}_{\text{FM}}+w_{R}\mathcal{L}_{R}+w_{H}\mathcal{L}_{H}$, where $w_{\text{FM}},w_{R},w_{H}$ are scalar weight. Performance is primarily evaluated by the final flow matching loss, which indicates how well the GFlowNet has learned the target distribution.
Method Validation Flow Loss Test Flow Loss
VRAdam 1.25 1.33
AdamW 2.41 3.6
SGD+Nesterov 2.71 2.61
RMSProp 25.0 25.0
Table 3: GFlowNets results of optimizers on GridWorld reported through flow match loss.
5.4 Hyperparameter Sensitivity of Diffusion Models for Generative Modeling
Diffusion models have emerged as powerful generative models, particularly for high-fidelity image and video synthesis. However, their training can be computationally intensive and sensitive to hyperparameter choices. This benchmark assesses the robustness of VRAdam and the ease of finding effective hyperparameter settings when training a diffusion model. The dataset consists of $28\times 28$ grayscale MNIST images.
The diffusion trading process involves a forward noising process and a learned reverse denoising process. The reverse process, which generates images by iteratively removing noise, is powered by a UNet architecture with self-attention. This U-Net is specifically designed for diffusion tasks, incorporating time embeddings (derived from sinusoidal positional encodings processed by an MLP) that allow the network to condition its predictions on the current noise level (timestep). For these benchmarks, this U-Net due to its self-attention mechanism effectively captures global image context and long-range dependencies, which are vital for generating coherent images.
Figure 3: The mean squared error (MSE) training and validation loss over ten sweeps of hyperparameters using Bayesian optimization for a generative modeling task using a Transformer optimized with VRAdam and AdamW for [Hyperparameter in the appendix].
Crucially as seen in Fig. 3, we highlight the robustness of VRAdam to various hyperparameter choices during multiple sweeps by reporting the training and validation loss. VRdam demonstrates significantly more stable and robust performance than AdamW for training this diffusion model, with far smoother convergence. In contrast, AdamW exhibits widespread instability, with many hyperparameter settings leading to erratic training, failure to converge Karras et al. (2022), even as some specific choices lead to comparable performance to that of VRAdam . This robust performance across a wide range of its hyperparameters suggests that VRdam can achieve satisfactory results with minimal tuning, and potentially even with its default settings, thereby reducing the need for extensive hyperparameter searches. We also provide the details regarding the performance of SGD and RMSProp in the appendix.
@@ 6 Related Work
Follow-up work on Adam The Adam optimizer was introduced in 2014 by Kingma and Ba (2017) by combining the previously existing concepts of momentum and scaling a base LR for each parameter based on second-order moment estimates. The base learning rate, however, remains hard-coded (potentially chosen through a learning rate scheduler) throughout training. Since then, several modifications to Adam have been introduced, such as NAdam Dozat (2016), RAdam Liu et al. , Adabelief Zhuang et al. (2020), and in particular AdamW Loshchilov and Hutter , which reintroduces weight decay in its original intention. LARS You et al. (2017) and LAMB You et al. (2020) compute learning rates for layers individually. More recent optimization techniques include LION, an automatically discovered alternative to signed momentum Chen et al. (2023), Sophia Liu et al. (2023a), which uses estimated diagonal entries of the Hessian as a precondition, sharpness-aware minimization methods Foret et al. (2021), and a modified LR in Adam for reinforcement learning Ellis et al. (2024).
Understanding training dynamics, convergence analysis, and edge of stability Another line of work focuses on understanding the dynamics of the training of deep neural networks as well as derive convergence properties and guarantees for commonly used optimizers. Wang and Choromanska (2025) provides a recent survey over the later and Reddi et al. (2019) provide an explicit convex example for which Adam does not converge. A popularized framework for understanding training dynamics in the continuous training flow and infinite network width limit was introduced in Jacot et al. (2018), extended by finite width corrections in Huang and Yau (2020), and developed for graph neural networks in Du et al. (2019). Lastly, the role of the edge of stability regime offers an empirically view and was analyzed in Cohen et al. (2024); Arora et al. (2022); Cohen et al. (2022); Damian et al. (2022); Song and Yun (2023); Wang et al. (2022)
Symplectic Optimization This research area derives discrete-time optimization algorithms by discretizing continuous-time Hamiltonian or Lagrangian flows using symplectic integrators, which exactly preserve the underlying geometric (symplectic) structure of the dynamical system. This approach guarantees long-term stability, energy-preservation properties (or controlled energy dissipation in the dissipative case), and can provide valuable insights into existing optimizers. Notable work includes Betancourt et al. (2018); França et al. (2020); Maddison et al. (2018); Duruisseaux and Leok (2023); Yuan and Zhang (2023).
@@ 7 Discussion
Limitations We note that due to limited compute resources we were not able to benchmark the performance of VRAdam on model sizes beyond 56 million parameters. The edge of stability regime remains not fully understood and questions regarding generalization capabilities remain.
Summary of work Motivated by physical perspectives for complex optimization scenarios and stability conditions along with the adaptive edge of stability, we developed a new optimizer VRAdam based on quartic kinetic energy terms. We analyzed its performance at the adaptive edge of stability and benchmark it against several optimizers in particular AdamW on image classification, language modeling, a generative task using GFlowNets as well as diffusion modeling where we report improved performance and robustness to hyperparameters.
@@ Appendix A Bounding NRQCD
Starting with the NRQCD Lagrangian density in Euclidean space:
@@
[…]
@@ A.1 Analysis of Boundedness
Appendix B Hyperparameters, Datasets and Model Architectures
B.1 Image Classification

### v2

@@ 5 Benchmarks
We benchmark VRAdam against various optimizers on several datasets for various tasks with diverse architectures. These benchmarks include image classification using a CNN in the CIFAR-10 dataset Krizhevsky & Hinton (2009), language modeling with a Transformer architecture Vaswani et al. (2023) on the WikiText2 dataset Merity et al. (2016), and Generative Flow Networks (GFLowNets) Avidan et al. (2025), on a grid world for sequence generation tasks Chevalier-Boisvert et al. (2023), and a training a large language model (LLM) such as the Generative Pretrained Transformer (GPT) Brown et al. (2020). These benchmarks represent a broad variety of deep learning architectures and tasks, ranging from older techniques for image classification to much newer and actively emerging models and training techniques such as for GFLowNets. We focus on the comparison with AdamW since its exceeding popularity and consistent performance but also report the performance of stochastic gradient descent (SGD) with momentum Qian (1999), root mean square propagation (RMSProp) Ruder (2017) and RAdam Liu et al. (2020) for all tasks expect for the LLM benchmark. In particular, we demonstrate the excellent convergence properties as well as performance improvements in most benchmarks for VRAdam. Note that we run Bayesian optimization for hyperparameter sweeps for the image classification, and language modeling tasks, with several evaluations in close proximity to the found minimum, and the results are statistically significant. We report the optimal value from this optimization procedure, however, we include shaded error bars in the appendix. The Bayesian optimization for the hyperparameters was conducted with the objective to minimize validation loss, however, we also report test loss. Hyperparameters for the task involving GFlowNets were picked at random due compute constraints. All details regarding model training, setup and datasets along with error envelopes using multiple runs are provided in Appendices D and F.
LLM benchmark: In order to demonstrate the effectiveness of VRAdam for training large networks, such as large language models Brown et al. (2020), we present the validation loss results of VRAdam against that of AdamW when training a 124M parameter GPT-2 model on the FineWebEdu-10B dataset from scratch Penedo et al. (2024). We report the validation loss after around  57000 steps or 2 epochs as well as the average training time per epoch on a H100 GPU using hyperparameters found in the Appendix.
Method Training time (s) Validation Loss
AdamW 48549.56 3.511
VRAdam 48522.40 3.447
Table 1: Comparison of training time and validation loss for GPT-2 training.
Method WikiText-2 Loss CIFAR-10 Loss GridWorld Flow Matching Loss
Validation Test Validation Test Validation Test
VRAdam 5.99 6.00 0.476 0.469 1.25 1.33
AdamW 6.47 6.50 0.522 0.565 2.41 3.60
RAdam 7.511 7.554 2.300 4.005 1.407 2.290
SGD+Nesterov NaN NaN 0.625 0.620 2.71 2.61
RMSProp NaN NaN 0.801 0.813 25.0 25.0
Table 2: Comparison of optimizer performance across three tasks: language modeling on WikiText-2, image classification on CIFAR-10, and flow matching on GridWorld.
@@ 6 Related Work
Follow-up work on Adam The Adam optimizer was introduced in 2014 by Kingma & Ba (2017) by combining the previously existing concepts of momentum and scaling a base LR for each parameter based on second-order moment estimates. The base learning rate, however, remains hard-coded (potentially chosen through a learning rate scheduler) throughout training. Since then, several modifications to Adam have been introduced, such as NAdam Dozat (2016), RAdam Liu et al. (2020), Adabelief Zhuang et al. (2020), and in particular AdamW Loshchilov & Hutter (2017), which reintroduces weight decay in its original intention. LARS You et al. (2017) and LAMB You et al. (2020) compute learning rates for layers individually. More recent optimization techniques include LION, an automatically discovered alternative to signed momentum Chen et al. (2023), Sophia Liu et al. (2023), which uses estimated diagonal entries of the Hessian as a precondition, sharpness-aware minimization methods Foret et al. (2021), and a modified LR in Adam for reinforcement learning Ellis et al. (2024).
Understanding training dynamics, convergence analysis, and edge of stability Another line of work focuses on understanding the dynamics of the training of deep neural networks as well as derive convergence properties and guarantees for commonly used optimizers. Wang & Choromanska (2025) provides a recent survey over the later and Reddi et al. (2019) provide an explicit convex example for which Adam does not converge. A popularized framework for understanding training dynamics in the continuous training flow and infinite network width limit was introduced in Jacot et al. (2018), extended by finite width corrections in Huang & Yau (2020), and developed for graph neural networks in Du et al. (2019). Lastly, the role of the edge of stability regime offers an empirically view and was analyzed in Cohen et al. (2024); Arora et al. (2022); Cohen et al. (2022); Damian et al. (2022); Song & Yun (2023); Wang et al. (2022)
Symplectic Optimization This research area derives discrete‐time optimization algorithms by discretizing continuous‐time Hamiltonian or Lagrangian flows using symplectic integrators, which exactly preserve the underlying geometric (symplectic) structure of the dynamical system. This approach guarantees long‐term stability, energy‐preservation properties (or controlled energy dissipation in the dissipative case), and can provide valuable insights into existing optimizers. Notable work includes Betancourt et al. (2018); França et al. (2020); Maddison et al. (2018); Duruisseaux & Leok (2023); Yuan & Zhang (2023).
@@ 7 Discussion
Limitations We note that due to limited compute resources we were not able to benchmark the performance of VRAdam on model sizes beyond GPT 124M for a longer period. The edge of stability regime remains not fully understood and questions regarding generalization capabilities remain.
Summary of work Motivated by physical perspectives for complex optimization scenarios and stability conditions along with the adaptive edge of stability, we developed a new optimizer VRAdam based on quartic kinetic energy terms. We analyzed its performance at the adaptive edge of stability and benchmark it against several optimizers in particular AdamW on image classification, language modeling, and a generative task using GFlowNets where we report improved performance and robustness to hyperparameters.
@@ Appendix A Bounding NRQCD
By explicitly breaking certain symmetries—Lorentz invariance in NRQCD and time-translation in time crystals— higher-order kinetic terms paradoxically enhance stability through topological protection mechanisms and the generation of emergent length/time scales Niemi (2021); Guha & Ghose-Choudhury (2019).
As a demonstration of this phenomenon, we can consider Nonrelativistic QCD (NRQCD), which is an effective field theory that expands full quantum chromodynamics in inverse powers of a heavy‐quark mass $m$ Assi et al. (2023). The bilinear heavy‐quark Lagrangian $\mathcal{L}$ up to $O(1/m^{3})$ reads,
@@
$D_{0}$ $\rightarrow-iE\quad\text{(time evolution operator)}$ (22)
$\mathbf{D}$ $\rightarrow i\vec{p}\quad\text{(spatial covariant derivative)}$ (23)
$\mathbf{D}^{2}$ $\rightarrow-p^{2}\quad\text{(squared spatial derivative)}$ (24)
$\mathbf{D}^{4}$ $\rightarrow p^{4}\quad\text{(quartic spatial derivative)}$ (25)
@@ A.1 Analysis of Boundedness
Appendix B Physics analysis
In this section, we provide a stability analysis from starting with a physical perspective, starting from our quartic Lagrangian from Eq. 3, we derive the 1D dynamics which we discretize implicitly through the state‑dependent step
$\eta_{t}=\frac{\alpha_{0}}{1+\min(\beta_{3}|v_{t}|^{2},\alpha_{1})},\qquad v_{t}=\beta_{1}v_{t-1}+(1-\beta_{1})\nabla V(x_{t-1}),\qquad x_{t}=x_{t-1}-\eta_{t}v_{t}.$ (37)
For a quadratic potential $V(x)=\tfrac{1}{2}x^{\top}Hx$ with $0\prec H\preceq LI$, we work in the eigenbasis $H=Q^{\top}\mathrm{diag}(h_{i})Q$. Writing coordinates as $(x_{i},v_{i})$ and $h\in(0,L]$, the one‑dimensional update reads
$v=\beta_{1}u+(1-\beta_{1})h,x,\qquad x^{+}=x-\eta v,\qquad y:=\eta h,\qquad\alpha:=1-\beta_{1},$ (38)
where $u:=v_{t-1}$ and $\eta\in[\eta_{\min},\alpha_{0}]$ with $\eta_{\min}=\alpha_{0}/(1+\alpha_{1})$. We introduce an energy that mirrors the quartic kinetic term and couples consistently to the curvature,
$\mathcal{E}h(x,v)=\frac{1}{2}h,x^{2}+\sigma,xv+\frac{\gamma}{2h},v^{2}+\frac{\delta}{4},v^{4},$ (39)
and set
$\sigma=\zeta,\frac{\beta_{1}}{2},\eta_{\min},\qquad\gamma=\zeta,\frac{\eta_{\min}}{1-\beta_{1}},\qquad\delta\geq\frac{\alpha_{0}^{2}}{\beta_{3}},\qquad\zeta\in(0,1).$ (40)
A direct expansion of $\mathcal{E}h(x^{+},v)-\mathcal{E}h(x,u)$ gives a quadratic form in $(x,u)$ plus the quartic difference $(v^{4}-u^{4})/4$. Because the coefficients of the quadratic form are convex polynomials of $y=\eta h$, the worst case over $y\in[\eta_{\min}h,\alpha_{0}h]$ occurs at the endpoints. Evaluating there, and using the adaptive edge‑of‑stability bound for momentum
$\alpha_{0}L<\frac{2(1+\beta_{1})}{1-\beta_{1}},$ (41)
one obtains constants $c_{x},c_{u}>0$ (depending only on $\beta_{1},\alpha_{0},\eta_{\min}$) such that, uniformly for all $h\in[\mu,L]$ and all $\eta\in[\eta_{\min},\alpha_{0}]$,
$\mathcal{E}h(x^{+},v)-\mathcal{E}h(x,u)\leq-,c_{x},h,x^{2}-,c_{u},\frac{u^{2}}{h}+\frac{\delta}{4},(v^{4}-u^{4}).$ (42)
The gate implies $\eta\leq\alpha_{0}/(1+\beta_{3}v^{2})$ and hence $\tfrac{1}{2}\eta^{2}v^{2}\leq\alpha_{0}^{2}/(8\beta_{3})$. The choice $\delta\geq\alpha_{0}^{2}/\beta_{3}$ therefore ensures that whenever $|v|\geq 1/\sqrt{\beta_{3}}$ the quartic contribution dominates any potential increase from the discretization term and contributes strict dissipation; when $|v|\leq 1/\sqrt{\beta_{3}}$ the negative quadratic terms already control the step. Summing over coordinates yields a global energy
$\mathcal{E}(x,v)=\sum{i=1}^{d}\mathcal{E}{h_{i}}(x_{i},v_{i}),$ (43)
and constants $\kappa_{1},\kappa_{2},\kappa_{4}>0$ such that
$\mathcal{E}(x_{t},v_{t})-\mathcal{E}(x_{t-1},v_{t-1})\leq-\kappa_{1}x_{t-1}^{\top}Hx_{t-1}-\kappa_{2}v_{t-1}^{\top}H^{-1}v_{t-1}-\kappa_{4}\sum_{i=1}^{d}\max\left\{0,v_{t,i}^{4}-\frac{1}{\beta_{3}^{2}}\right\}.$ (44)
This inequality formalizes the physical picture in Figure 1. The first two terms express curvature‑weighted exchange between potential and quadratic kinetic energy with net dissipation; the quartic term is a brake that activates precisely in high‑velocity regimes created near the adaptive edge of stability. The same mechanism explains the reduction in ringing and the lower sharpness observed in Figure 2: when $|v_{t}|$ grows, the gate reduces $\eta_{t}$ and the quartic channel increases dissipation, pushing the dynamics back to a low‑velocity regime. Two immediate consequences follow from the construction. First, the instantaneous adaptive stability threshold increases with the measured velocity:
$L_{\mathrm{EoS}}(t)=\frac{2(1+\beta_{1})}{(1-\beta_{1})\eta_{t}}=\frac{2(1+\beta_{1})}{(1-\beta_{1})\alpha_{0}}\left(1+\min(\beta_{3}|v_{t}|^{2},\alpha_{1})\right),$ (45)
so the method moves away from instability as oscillations grow. Second, each parameter update is uniformly bounded in norm by the gate,
$|x_{t}-x_{t-1}|=\eta_{t}|v_{t}|=\frac{\alpha_{0}|v_{t}|}{1+\beta_{3}|v_{t}|^{2}}\leq\frac{\alpha_{0}}{2\sqrt{\beta_{3}}},$ (46)
which prevents runaway steps and is not available to classical momentum. These properties are consistent with the design of Algorithm 1 and the empirical behavior reported in the analysis section.
Appendix C Convergence Proof for VRAdam
The proof is based on the work of the proof in Défossez et al. (2022).
There, the authors derive in equation A.37 the bound
$\underbrace{\tfrac{1}{2R}\sum_{n=1}^{N}\tfrac{\alpha_{n}}{\Omega_{n}}\sum_{k=0}^{n-1}\beta_{1}^{k}\,\mathbb{E}\!\left[\|G_{n-k}\|_{2}^{2}\right]}_{\mathbf{A}}$ $\;\leq\;F(x_{0})-F_{*}$
$\quad+\;\underbrace{\alpha_{N}^{2}L^{2}\sum_{n=1}^{N}\mathbb{E}\!\left[\|u_{n}\|_{2}^{2}\right]}_{\mathbf{B}}$
$\quad+\;\underbrace{\tfrac{\alpha_{N}^{3}L^{2}}{4R\sqrt{1-\beta_{1}}}\sum_{n=1}^{N}\sum_{l=1}^{n-1}\mathbb{E}\!\left[\|u_{n-l}\|_{2}^{2}\right]\sum_{k=l}^{n-1}\tfrac{\beta_{1}^{k}}{\sqrt{k}}}_{\mathbf{C}}$
$\quad+\;\underbrace{3\alpha_{N}R\sqrt{1-\beta_{1}}\sum_{n=1}^{N}\sum_{k=0}^{n-1}\!\Big(\tfrac{\beta_{1}}{\beta_{2}}\Big)^{k}\!\sqrt{k+1}\;\mathbb{E}\!\left[\|U_{n-k}\|_{2}^{2}\right]}_{\mathbf{D}}.$
By bounding $\alpha_{n}$ with $\alpha_{\min}$ in expression $A$, and $\alpha_{N}$ with $\alpha_{\max}$} in expressions $B,C,D$ we obtain the final result.
Appendix D Hyperparameters, Datasets and Model Architectures
Note, that we recommend $\beta_{3}=1.0$ as the default setting for VRAdam.
D.1 Image Classification